OpenStudy (anonymous):

solve((x+3)/(x^2-4)-1/(x+2)=1/(x-2), x)

OpenStudy (anonymous):

multiply both sides by (x^2-4) (x+3) - (x-2) =(x+2) x+3-x+2-x-2 = 0 -x+3 =0 x=3 ----> solution